Determine and interpret Pearson’s correlation coefficient.
1
Expert's answer
2019-07-11T13:14:11-0400

The experimental data set could have some correlations in the data. These correlations between sets of data is commonly measured by the Pearson Correlation Coefficient.

For example. is these are two data sets: A and B, and we want to know whether these values are correlated. For this purpose, scientists calculate Pearson Correlation Coefficient, and then, by analyzing the value of this coefficient, they could say whether there is correlation between data A and B, or not.


Pearson Correlation Coefficient for two data sets X and Y is calculated as follows:


"r_{xy}=\frac{\sum_{i=1}^{n}(x_i-\bar{x})(y_i-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_i-\bar{x})^2}\sqrt{\sum_{i=1}^{n}(y_i-\bar{y})^2}}"